what is the average rate of change of the function from x=3 to x=5 using this function: f(x)=2-xsquared
Answer by RAY100(1637)

f(x) = 2- x^2
.
.
f(3) = 2-3^2 = 2-9= -7
.
f(5) =2-5^2 = 2-25 = -23
.
(3, -7) ans (5, -23)
.
.
rate of change is slope,,,m= (y2-y1) /(x2-x1)
.
m= (-23-(-7)) / 5-3 = -16 /2 = -8
.
average rate of change of function is -8
.